OpenStudy (anonymous):

Someone please help me im having serious trouble its frustrating and im in desperate need i really need someone that can walk me through these problems please

OpenStudy (anonymous):

just please if you're willing to help me through a couple of steps u will get every medal

OpenStudy (anonymous):

@UsukiDoll are u able to help me

OpenStudy (anonymous):

i really need it i mean really i'm a freshman nd if i don't get this done by tomorrow im in serious trouble so please just please

OpenStudy (anonymous):

i will give you a medal in everything you answer @UsukiDoll

OpenStudy (usukidoll):

what is it

OpenStudy (anonymous):

Indicate which property is illustrated in Step 5. Step 1 6x + 7 + 7x + 6 = (6x + 7) + (7x + 6) Step 2 = 6x + (7 + 7x) + 6 Step 3 = 6x + (7x + 7) + 6 Step 4 = (6x + 7x) + (7 + 6) Step 5 = (6 + 7)x + (7 + 6) Step 6 = 13x + 13 @UsukiDoll

OpenStudy (usukidoll):

O_O oh darn

OpenStudy (anonymous):

Distributive property, I think.

OpenStudy (usukidoll):

it's either associative property or cumulative property or distributive property not sure

zepdrix (zepdrix):

So it's .. factoring which ya maybe is called the Distributive Property (in reverse) or something..

OpenStudy (usukidoll):

what those two said...sorry that stuff was eons ago and may never be used again .

OpenStudy (anonymous):

Indicate which property is illustrated in Step 6. Step 1 9x + 9 + 11x + 9 = (9x + 9) + (11x + 9) Step 2 = 9x + (9 + 11x) + 9 Step 3 = 9x + (11x + 9) + 9 Step 4 = (9x + 11x) + (9 + 9) Step 5 = (9 + 11)x + (9 + 9) Step 6 = 20x + 18

OpenStudy (anonymous):

There is probably a property that says you can apply properties in reverse. They are usually rather thorough. Actually, I think the symmetric property pretty much does say that.

OpenStudy (coconutjj):

distributive property

OpenStudy (anonymous):

Anyway, that one is the "uh, addition actually works" property. Additive property of equality? Or something.

OpenStudy (anonymous):

arithmetic fact

OpenStudy (anonymous):

There is some property that says "a = a" and the proof for it is probably 50 pages long and involves calculus.

OpenStudy (anonymous):

Indicate which property is illustrated in Step 6. Step 1 5 - 8x - 5 + 9x = (5 - 8x) + (-5 + 9x) Step 2 = (5 - 8x) + (9x - 5) Step 3 = 5 + (-8x + 9x) - 5 Step 4 = 5 + (-8 + 9)x - 5 Step 5 = 5 + x - 5 Step 6 = x + 5 - 5 Step 7 = x + 0 Step 8 = x

OpenStudy (anonymous):

Additive inverse property, I think.

OpenStudy (anonymous):

commutative distributive identity associative

OpenStudy (anonymous):

This is one of the ugliest web pages I have ever seen but it does have a nice list of properties: http://www.coolmath.com/prealgebra/06-properties/

OpenStudy (anonymous):

Oh, commutative. I was looking at the wrong part. They're probably showing that 5 + x = x + 5.

OpenStudy (anonymous):

Indicate which property is illustrated in Step 5. Step 1 4x + 9 + 5x + 9 = (4x + 9) + (5x + 9) Step 2 = 4x + (9 + 5x) + 9 Step 3 = 4x + (5x + 9) + 9 Step 4 = (4x + 5x) + (9 + 9) Step 5 = (4 + 5)x + (9 + 9) Step 6 = 9x + 18

OpenStudy (anonymous):

We already did this one! You should know it.

OpenStudy (anonymous):

@cassytaylor92 youre back

OpenStudy (anonymous):

uhmm , yes :) @Starr_Dynasty

OpenStudy (wolf1728):

SACAPUNTUS - I believe that 50 page "proof" for something like 1+0 =1 is from the 3 volume Principia Mathematica by Bertrand Russell.

OpenStudy (anonymous):

lol well thts gr8 @cassytaylor92